OpenStudy (anonymous):

Three brothers spent different amounts of money on their recent vacation. Matt spent $25 less than Todd. Jeff spent 3 times as much as Matt. How much did each brother spend if they spent a total of $275?

OpenStudy (anonymous):

x + x-25 + 3(x-25) = 275 5x -100 = 275 5x = 375 x = 75

OpenStudy (anonymous):

I thought that is how you set up the problem, but do you multiply 3 by (x-25) first?

OpenStudy (anonymous):

Todd is x Matt is x- 25 Jeff is 3(x-25)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

yes I understand that part

OpenStudy (anonymous):

x = 75 so you just plug that in

OpenStudy (anonymous):

Okay. How did you get 5x?

OpenStudy (anonymous):

275 + 100 = 375 then we have to add up all the exponents. so we have 3x + x + x = 5x

OpenStudy (anonymous):

in order ti isolate x we must divide 5x/ 375 so x =